OpenStudy (anonymous):

Create an addition word problem and a subtraction word problem using fractions. Show how to find the solution to each problem. Once you find the solutions to your problems, rewrite the solutions as either a terminating or nonterminating decimal.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Hero

OpenStudy (phi):

Create an addition word problem can you come up with a problem with you and your friend both ate part of a pie. How much did you and your friend eat all together ?

OpenStudy (anonymous):

ok so there are 8 pieces of pie and i ate 1/2 and he ate 1/3 so we ate 5/6

OpenStudy (phi):

yes, that is ok now change 1/2 to a decimal change 1/3 to a decimal

OpenStudy (anonymous):

1/2=0.5 1/3=0.33333333333333333333333333333333

OpenStudy (phi):

ok. btw I would not say 8 pieces of 8. Just say you have a whole cake 1/2+ 1/3 = 0.833.... if you change 5/6 to a decimal you should get 0.8333...

OpenStudy (anonymous):

ok